# Coarse-Grained Games: A Framework for Bounded Perception in Game Theory

### v1

@@ 4.3.3 Loss of Competitiveness in the Game
$\forall s_{l},s_{-l},s^{\prime}_{l},s^{\prime}_{-l},\quad\psi_{\text{EMP}}(\phi_{k}(u_{l}(s_{l},s_{-l})))=\psi_{\text{EMP}}(\phi_{k}(u_{l}(s^{\prime}_{l},s^{\prime}_{-l}))).$
@@
In this coarse-grained game, the optimal response for player 1 is to choose “Cooperation” with probability $\frac{3}{13}$ and “Defect” with probability $\frac{10}{13}$. However, for player 2, all payoffs have been mapped to the same numerical value under the EMP transformation, meaning that their expected utility remains constant regardless of their strategy selection. As a result, player 2 faces an equilibrium selection problem.
@@ 5.1 Incidental Gain-Loss Differential
Determining whether an incidental advantage or disadvantage occurs from the subjective perspective, and if so, under what conditions, can be efficiently analyzed using the standard procedure for finding Nash equilibria in games. This is because the advantage in a coarse-grained game for player $k$ is represented by the difference between:
@@
In a coarse-grained game with two players, we determine each player’s perceived Nash equilibrium based on their respective payoff matrices:
@@
$\psi_{\text{EMP}}(\phi_{1}(u_{1}(s^{{\dagger}_{1}}_{1},s^{{\dagger}_{1}}_{2})))\geq\psi_{\text{EMP}}(\phi_{1}(u_{1}(s_{j_{1}},s^{{\dagger}_{1}}_{2}))),$
@@
$\psi_{\text{EMP}}(\phi_{1}(u_{2}(s^{{\dagger}_{1}}_{1},s^{{\dagger}_{1}}_{2})))\geq\psi_{\text{EMP}}(\phi_{1}(u_{2}(s^{{\dagger}_{1}}_{1},s_{j_{2}}))),$
@@
$\psi_{\text{EMP}}(\phi_{2}(u_{1}(s^{{\dagger}_{2}}_{1},s^{{\dagger}_{2}}_{2})))\geq\psi_{\text{EMP}}(\phi_{2}(u_{1}(s_{j_{1}},s^{{\dagger}_{2}}_{2}))),$
@@
$\psi_{\text{EMP}}(\phi_{2}(u_{2}(s^{{\dagger}_{2}}_{1},s^{{\dagger}_{2}}_{2})))\geq\psi_{\text{EMP}}(\phi_{2}(u_{2}(s^{{\dagger}_{2}}_{1},s_{j_{2}}))),$
@@
$\Delta^{M^{\prime}_{1}}_{1}=\psi_{\text{EMP}}(\phi_{1}(u_{1}(s^{{\dagger}_{1}}_{1},s^{{\dagger}_{2}}_{2})))-\psi_{\text{EMP}}(\phi_{1}(u_{1}(s^{{\dagger}_{1}}_{1},s^{{\dagger}_{1}}_{2})))$
@@
In a coarse-grained game, each player $k$ perceives a subjective, coarse-grained payoff matrix $M^{\prime}_{k}$. The perceived Nash equilibrium strategies for player $k$ are denoted as $(s^{{\dagger}_{k}}_{1},\dots,s^{{\dagger}_{k}}_{n})$, satisfying:
$\psi_{\text{EMP}}(\phi_{k}(u_{k}(s^{{\dagger}_{k}}_{1},\dots,s^{{\dagger}_{k}}_{k},\dots,s^{{\dagger}_{k}}_{n})))\geq\psi_{\text{EMP}}(\phi_{k}(u_{k}(s^{{\dagger}_{k}}_{1},\dots,s_{j_{k}},\dots,s^{{\dagger}_{k}}_{n}))),$
@@
$\Delta^{M^{\prime}_{k}}_{k}=\psi_{\text{EMP}}(\phi_{k}(u_{k}(s^{{\dagger}_{1}}_{1},\dots,s^{{\dagger}_{k}}_{k},\dots,s^{{\dagger}_{n}}_{n})))-\psi_{\text{EMP}}(\phi_{k}(u_{k}(s^{{\dagger}_{k}}_{1},\dots,s^{{\dagger}_{k}}_{k},\dots,s^{{\dagger}_{k}}_{n}))).$
@@
However, in the coarse-grained game, each payoff is transformed by $\phi_{k}$ and $\psi_{k}$ (in this paper, $\psi_{\text{EMP}}$ for all players), requiring the previous equations for indifference to be modified as follows.
@@
$p_{1_{2}}\cdot\psi_{\text{EMP}}(\phi_{1}(u_{1}(s_{1_{1}},s_{1_{2}})))+p_{2_{2}}\cdot\psi_{\text{EMP}}(\phi_{1}(u_{1}(s_{1_{1}},s_{2_{2}})))$
$\quad=p_{1_{2}}\cdot\psi_{\text{EMP}}(\phi_{1}(u_{1}(s_{2_{1}},s_{1_{2}})))+p_{2_{2}}\cdot\psi_{\text{EMP}}(\phi_{1}(u_{1}(s_{2_{1}},s_{2_{2}}))),$
$p_{1_{1}}\cdot\psi_{\text{EMP}}(\phi_{1}(u_{2}(s_{1_{1}},s_{1_{2}})))+p_{2_{1}}\cdot\psi_{\text{EMP}}(\phi_{1}(u_{2}(s_{2_{1}},s_{1_{2}})))$
$\quad=p_{1_{1}}\cdot\psi_{\text{EMP}}(\phi_{1}(u_{2}(s_{1_{1}},s_{2_{2}})))+p_{2_{1}}\cdot\psi_{\text{EMP}}(\phi_{1}(u_{2}(s_{2_{1}},s_{2_{2}}))).$
@@
$p_{1_{2}}\cdot\psi_{\text{EMP}}(\phi_{2}(u_{1}(s_{1_{1}},s_{1_{2}})))+p_{2_{2}}\cdot\psi_{\text{EMP}}(\phi_{2}(u_{1}(s_{1_{1}},s_{2_{2}})))$
$\quad=p_{1_{2}}\cdot\psi_{\text{EMP}}(\phi_{2}(u_{1}(s_{2_{1}},s_{1_{2}})))+p_{2_{2}}\cdot\psi_{\text{EMP}}(\phi_{2}(u_{1}(s_{2_{1}},s_{2_{2}}))),$
$p_{1_{1}}\cdot\psi_{\text{EMP}}(\phi_{2}(u_{2}(s_{1_{1}},s_{1_{2}})))+p_{2_{1}}\cdot\psi_{\text{EMP}}(\phi_{2}(u_{2}(s_{2_{1}},s_{1_{2}})))$
$\quad=p_{1_{1}}\cdot\psi_{\text{EMP}}(\phi_{2}(u_{2}(s_{1_{1}},s_{2_{2}})))+p_{2_{1}}\cdot\psi_{\text{EMP}}(\phi_{2}(u_{2}(s_{2_{1}},s_{2_{2}}))).$
@@
Here, $\Phi_{1}$ represents the application of $\phi_{1}$ to each outcome of $u_{1}(s_{j_{1}},s_{j_{2}})$, while $\Psi_{\text{EMP}}$ is a function that further applies $\psi_{\text{EMP}}$ to this result. In the current case, the calculation formula for each strategy profile is given by:
$p_{j_{1}}\cdot p_{j_{2}}\cdot\psi_{\text{EMP}}(\phi_{1}(u_{1}(s_{j_{1}},s_{j_{2}}))).$
@@
Applying the notation defined for the two-player case, the expected payoff of player $k$ at Nash equilibrium in the coarse-grained game in $M^{\prime}_{k}$ is given by the following equation:
@@ 6.2.1 Fulfillment of the Folk Theorem in Coarse-Grained Matrix
$\psi_{\text{EMP}}(\phi_{k}(u_{k}(s_{j_{1}},\dots,s_{j_{n}})))$
where $u_{k}(s_{j_{1}},\dots,s_{j_{n}})$ is the true payoff in the base matrix, $\phi_{k}$ is the coarse-graining function mapping real payoffs into a partitioned structure, and $\psi_{\text{EMP}}$ is the EMP, which assigns a representative numerical value to the coarse-grained partition.
Since the coarse-graining function $\phi_{k}$ partitions payoffs into intervals or sets, and EMP assigns a single representative value per partition, the resulting payoff remains a well-defined real number. This ensures that subjectively, players perceive the game as having well-defined payoffs.
@@
Step 2: Subjective Observability of Opponent Strategies. In a standard perfect monitoring setup, players can observe opponents’ strategies exactly. However, in a coarse-grained game, player $k$ only perceives payoffs through $\phi_{k}$, meaning they do not see precisely which strategy was played but instead recognize the outcome at a coarser level.
@@ 6.2.2 Intersubjective Discrepancy in Perceived Sufficient Discount Rates
of Lemma 3 We prove that in a coarse-grained game, the sufficient discount factor for cooperation may differ depending on the observer’s perspective.
@@
However, in a coarse-grained game, the perceived utility of player $l$ from the perspective of player $k$ is given by:
$V^{M^{\prime}_{k}}_{l}=\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{k}}_{l})^{t}\psi_{k}(\phi_{k}(u_{l}(s_{t}))),$
@@
$\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{k}}_{l})^{t}\psi_{\text{EMP}}(\phi_{k}(T_{C}))$ $\geq\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{k}}_{l})^{t}\psi_{\text{EMP}}(\phi_{k}(u_{l}(s^{\prime}_{t}))).$
@@
$\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{l}}_{l})^{t}\psi_{\text{EMP}}(\phi_{l}(T_{C}))$ $\geq\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{l}}_{l})^{t}\psi_{\text{EMP}}(\phi_{l}(u_{l}(s^{\prime}_{t}))).$
@@
$\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{k}}_{l})^{t}\psi_{\text{EMP}}(\phi_{k}(T_{C}))$ $\geq\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{k}}_{l})^{t}\psi_{\text{EMP}}(\phi_{k}(u_{l}(s^{\prime}_{t})))$
$\iff\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{k}}_{l})^{t}\psi_{\text{EMP}}(\phi_{l}(T_{C}))\geq\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{k}}_{l})^{t}\psi_{\text{EMP}}(\phi_{l}(u_{l}(s^{\prime}_{t}))).$
However, this equivalence does not always hold. Given the definitions of $\psi_{\text{EMP}}$, $\phi_{k}$, and $\phi_{l}$, there is no guarantee that the value of $u_{l}(s_{t})$ is preserved under coarse-graining. In particular, $\phi_{k}$ and $\phi_{l}$ may map $u_{l}(s_{t})$ to different coarse-grained values, leading to discrepancies in the perceived utilities.
@@ 6.3 Failure of Cooperation
$\exists G\in\mathfrak{G}_{k},\quad\forall j_{k},\quad\psi_{\text{EMS}}(\phi_{k}(u_{k}(s_{j_{1}},s_{j_{2}},\dots,s_{j_{k}},\dots,s_{j_{n}})))\in G.$
@@ 7.1 Minor Model Change
In this paper, we aim to demonstrate the applicability of coarse-grained games in social sciences by exploring several modeling approaches in economics. The first analysis focuses on how consumers’ coarse-grained perception can influence a firm’s decision to implement minor model changes in its products. We formulate the following game:
@@
In this case, an equilibrium selection problem arises. Whether $C$ decides to buy $m_{2}$ depends on external factors, such as the consumer’s preference for new products or the level of trust they place in the dealer’s recommendation.
@@
In this coarse-grained payoff matrix, despite the difference in resolution between the dealer and the consumer in perceiving product distinctions, they ultimately reach the same conclusion. The dealer sells $m_{3}$, and the consumer $C$ purchases it, making this the unique Nash equilibrium.
@@
    Equilibrium Selection Problem in Minor Model Changes: Minor model changes can lead to an equilibrium selection problem unless the improvement in product quality results in a perceived utility gain that exceeds the consumer’s coarse-grained perception threshold. Without such a gain, the consumer’s choice between the old and new models depends on the focal point rather than a direct evaluation of quality differences. Within a coarse-grained game framework, an insufficiently significant model change—even if it objectively improves product quality—does not lead to competition based on quality differences. Instead, the consumer’s preference for novelty, the dealer’s persuasion skills, and the effectiveness of advertising become the dominant factors influencing the decision.

### v2

@@ 4.3.3 Loss of Competitiveness in the Game
$\forall s_{l},s_{-l},s^{\prime}_{l},s^{\prime}_{-l},\quad\psi_{\text{EMP}}(\varphi_{k}(u_{l}(s_{l},s_{-l})))=\psi_{\text{EMP}}(\varphi_{k}(u_{l}(s^{\prime}_{l},s^{\prime}_{-l}))).$
@@
In this CGG, the optimal response for player 1 is to choose “Cooperation” with probability $\frac{3}{13}$ and “Defect” with probability $\frac{10}{13}$. However, for player 2, all payoffs have been mapped to the same numerical value under the EMP transformation, meaning that their expected utility remains constant regardless of their strategy selection. As a result, player 2 faces an equilibrium selection problem.
@@ 5.1 Incidental Gain-Loss Differential
Determining whether an incidental advantage or disadvantage occurs from the subjective perspective, and if so, under what conditions, can be efficiently analyzed using the standard procedure for finding Nash equilibria in games. This is because the advantage in a CGG for player $k$ is represented by the difference between:
@@
In a CGG with two players, we determine each player’s perceived Nash equilibrium based on their respective payoff matrices:
@@
$\psi_{\text{EMP}}(\varphi_{1}(u_{1}(s^{{\dagger}_{1}}_{1},s^{{\dagger}_{1}}_{2})))\geq\psi_{\text{EMP}}(\varphi_{1}(u_{1}(s_{j_{1}},s^{{\dagger}_{1}}_{2}))),$
@@
$\psi_{\text{EMP}}(\varphi_{1}(u_{2}(s^{{\dagger}_{1}}_{1},s^{{\dagger}_{1}}_{2})))\geq\psi_{\text{EMP}}(\varphi_{1}(u_{2}(s^{{\dagger}_{1}}_{1},s_{j_{2}}))),$
@@
$\psi_{\text{EMP}}(\varphi_{2}(u_{1}(s^{{\dagger}_{2}}_{1},s^{{\dagger}_{2}}_{2})))\geq\psi_{\text{EMP}}(\varphi_{2}(u_{1}(s_{j_{1}},s^{{\dagger}_{2}}_{2}))),$
@@
$\psi_{\text{EMP}}(\varphi_{2}(u_{2}(s^{{\dagger}_{2}}_{1},s^{{\dagger}_{2}}_{2})))\geq\psi_{\text{EMP}}(\varphi_{2}(u_{2}(s^{{\dagger}_{2}}_{1},s_{j_{2}}))),$
@@
$\Delta^{M^{\prime}_{1}}_{1}=\psi_{\text{EMP}}(\varphi_{1}(u_{1}(s^{{\dagger}_{1}}_{1},s^{{\dagger}_{2}}_{2})))-\psi_{\text{EMP}}(\varphi_{1}(u_{1}(s^{{\dagger}_{1}}_{1},s^{{\dagger}_{1}}_{2})))$
@@
In a CGG, each player $k$ perceives a subjective, coarse-grained payoff matrix $M^{\prime}_{k}$. The perceived Nash equilibrium strategies for player $k$ are denoted as $(s^{{\dagger}_{k}}_{1},\dots,s^{{\dagger}_{k}}_{n})$, satisfying:
$\psi_{\text{EMP}}(\varphi_{k}(u_{k}(s^{{\dagger}_{k}}_{1},\dots,s^{{\dagger}_{k}}_{k},\dots,s^{{\dagger}_{k}}_{n})))\geq\psi_{\text{EMP}}(\varphi_{k}(u_{k}(s^{{\dagger}_{k}}_{1},\dots,s_{j_{k}},\dots,s^{{\dagger}_{k}}_{n}))),$
@@
$\Delta^{M^{\prime}_{k}}_{k}=\psi_{\text{EMP}}(\varphi_{k}(u_{k}(s^{{\dagger}_{1}}_{1},\dots,s^{{\dagger}_{k}}_{k},\dots,s^{{\dagger}_{n}}_{n})))-\psi_{\text{EMP}}(\varphi_{k}(u_{k}(s^{{\dagger}_{k}}_{1},\dots,s^{{\dagger}_{k}}_{k},\dots,s^{{\dagger}_{k}}_{n}))).$
@@
However, in the CGG, each payoff is transformed by $\varphi_{k}$ and $\psi_{k}$ (in this paper, $\psi_{\text{EMP}}$ for all players), requiring the previous equations for indifference to be modified as follows.
@@
$p_{1_{2}}\cdot\psi_{\text{EMP}}(\varphi_{1}(u_{1}(s_{1_{1}},s_{1_{2}})))+p_{2_{2}}\cdot\psi_{\text{EMP}}(\varphi_{1}(u_{1}(s_{1_{1}},s_{2_{2}})))$
$\quad=p_{1_{2}}\cdot\psi_{\text{EMP}}(\varphi_{1}(u_{1}(s_{2_{1}},s_{1_{2}})))+p_{2_{2}}\cdot\psi_{\text{EMP}}(\varphi_{1}(u_{1}(s_{2_{1}},s_{2_{2}}))),$
$p_{1_{1}}\cdot\psi_{\text{EMP}}(\varphi_{1}(u_{2}(s_{1_{1}},s_{1_{2}})))+p_{2_{1}}\cdot\psi_{\text{EMP}}(\varphi_{1}(u_{2}(s_{2_{1}},s_{1_{2}})))$
$\quad=p_{1_{1}}\cdot\psi_{\text{EMP}}(\varphi_{1}(u_{2}(s_{1_{1}},s_{2_{2}})))+p_{2_{1}}\cdot\psi_{\text{EMP}}(\varphi_{1}(u_{2}(s_{2_{1}},s_{2_{2}}))).$
@@
$p_{1_{2}}\cdot\psi_{\text{EMP}}(\varphi_{2}(u_{1}(s_{1_{1}},s_{1_{2}})))+p_{2_{2}}\cdot\psi_{\text{EMP}}(\varphi_{2}(u_{1}(s_{1_{1}},s_{2_{2}})))$
$\quad=p_{1_{2}}\cdot\psi_{\text{EMP}}(\varphi_{2}(u_{1}(s_{2_{1}},s_{1_{2}})))+p_{2_{2}}\cdot\psi_{\text{EMP}}(\varphi_{2}(u_{1}(s_{2_{1}},s_{2_{2}}))),$
$p_{1_{1}}\cdot\psi_{\text{EMP}}(\varphi_{2}(u_{2}(s_{1_{1}},s_{1_{2}})))+p_{2_{1}}\cdot\psi_{\text{EMP}}(\varphi_{2}(u_{2}(s_{2_{1}},s_{1_{2}})))$
$\quad=p_{1_{1}}\cdot\psi_{\text{EMP}}(\varphi_{2}(u_{2}(s_{1_{1}},s_{2_{2}})))+p_{2_{1}}\cdot\psi_{\text{EMP}}(\varphi_{2}(u_{2}(s_{2_{1}},s_{2_{2}}))).$
@@
Here, $\Phi_{1}$ represents the application of $\varphi_{1}$ to each outcome of $u_{1}(s_{j_{1}},s_{j_{2}})$, while $\Psi_{\text{EMP}}$ is a function that further applies $\psi_{\text{EMP}}$ to this result. In the current case, the calculation formula for each strategy profile is given by:
$p_{j_{1}}\cdot p_{j_{2}}\cdot\psi_{\text{EMP}}(\varphi_{1}(u_{1}(s_{j_{1}},s_{j_{2}}))).$
@@
Applying the notation defined for the two-player case, the expected payoff of player $k$ at Nash equilibrium in the CGG in $M^{\prime}_{k}$ is given by the following equation:
@@ 6.2.1 Fulfillment of the Folk Theorem in Coarse-Grained Matrix
$\psi_{\text{EMP}}(\varphi_{k}(u_{k}(s_{j_{1}},\dots,s_{j_{n}})))$
where $u_{k}(s_{j_{1}},\dots,s_{j_{n}})$ is the true payoff in the base matrix, $\varphi_{k}$ is the coarse-graining function mapping real payoffs into a partitioned structure, and $\psi_{\text{EMP}}$ is the EMP, which assigns a representative numerical value to the coarse-grained partition.
Since the coarse-graining function $\varphi_{k}$ partitions payoffs into intervals or sets, and EMP assigns a single representative value per partition, the resulting payoff remains a well-defined real number. This ensures that subjectively, players perceive the game as having well-defined payoffs.
@@
Step 2: Subjective Observability of Opponent Strategies. In a standard perfect monitoring setup, players can observe opponents’ strategies exactly. However, in a CGG, player $k$ only perceives payoffs through $\varphi_{k}$, meaning they do not see precisely which strategy was played but instead recognize the outcome at a coarser level.
@@ 6.2.2 Intersubjective Discrepancy in Perceived Sufficient Discount Rates
of Lemma 3 We prove that in a CGG, the sufficient discount factor for cooperation may differ depending on the observer’s perspective.
@@
However, in a CGG, the perceived utility of player $l$ from the perspective of player $k$ is given by:
$V^{M^{\prime}_{k}}_{l}=\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{k}}_{l})^{t}\psi_{k}(\varphi_{k}(u_{l}(s_{t}))),$
@@
$\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{k}}_{l})^{t}\psi_{\text{EMP}}(\varphi_{k}(T_{C}))$ $\geq\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{k}}_{l})^{t}\psi_{\text{EMP}}(\varphi_{k}(u_{l}(s^{\prime}_{t}))).$
@@
$\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{l}}_{l})^{t}\psi_{\text{EMP}}(\varphi_{l}(T_{C}))$ $\geq\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{l}}_{l})^{t}\psi_{\text{EMP}}(\varphi_{l}(u_{l}(s^{\prime}_{t}))).$
@@
$\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{k}}_{l})^{t}\psi_{\text{EMP}}(\varphi_{k}(T_{C}))$ $\geq\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{k}}_{l})^{t}\psi_{\text{EMP}}(\varphi_{k}(u_{l}(s^{\prime}_{t})))$
$\iff\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{k}}_{l})^{t}\psi_{\text{EMP}}(\varphi_{l}(T_{C}))\geq\sum_{t=0}^{\infty}(\delta^{M^{\prime}_{k}}_{l})^{t}\psi_{\text{EMP}}(\varphi_{l}(u_{l}(s^{\prime}_{t}))).$
However, this equivalence does not always hold. Given the definitions of $\psi_{\text{EMP}}$, $\varphi_{k}$, and $\varphi_{l}$, there is no guarantee that the value of $u_{l}(s_{t})$ is preserved under coarse-graining. In particular, $\varphi_{k}$ and $\varphi_{l}$ may map $u_{l}(s_{t})$ to different coarse-grained values, leading to discrepancies in the perceived utilities.
@@ 6.3 Failure of Cooperation
$\exists G\in\mathfrak{G}_{k},\quad\forall j_{k},\quad\psi_{\text{EMS}}(\varphi_{k}(u_{k}(s_{j_{1}},s_{j_{2}},\dots,s_{j_{k}},\dots,s_{j_{n}})))\in G.$
@@ 7.1 Minor Model Change
In this paper, we aim to demonstrate the applicability of CGGs in social sciences by exploring several modeling approaches in economics. The first analysis focuses on how consumers’ coarse-grained perception can influence a firm’s decision to implement minor model changes in its products. We formulate the following game:
@@
In this case, an equilibrium selection problem arises. Whether $C$ decides to buy $m_{2}$ depends on external factors, such as the consumer’s preference for new products or their level of trust in the dealer’s recommendation.
@@
In this coarse-grained payoff matrix, despite the difference in resolution between the dealer and the consumer in perceiving product distinctions, they ultimately reach the same conclusion. Formally, there exist three pure-strategy Nash equilibria corresponding to the diagonal entries of the payoff matrix. However, the strategy profile (Buy $(\text{Buy }m_{3},\text{Sell }m_{3})$ Pareto-dominates the others, yielding strictly higher payoffs for both players. Hence, it stands out as the unique equilibrium under the assumption of payoff-maximizing rationality and mutual optimality selection. Although Pareto dominance may serve as a basis for focal point selection, the equilibrium in this context is driven by payoff improvements resulting from a minor model change that enhances product quality. As such, it differs fundamentally from focal point coordination in settings where such quality improvements are not perceived by the players.
@@
    Equilibrium Selection Problem in Minor Model Changes: Minor model changes can give rise to an equilibrium selection problem unless the associated improvement in product quality yields a perceived utility gain that exceeds the consumer’s coarse-grained perception threshold. When this threshold is not surpassed, consumers are unable to distinguish between the old and new models based on quality, and their choices are instead guided by focal points unrelated to product merit—such as a preference for novelty, the dealer’s persuasive tactics, or the influence of advertising. Conversely, when the perceived gain exceeds the threshold, the quality improvement itself functions as a natural focal point, leading to equilibrium selection driven by product merit rather than extrinsic cues.